Which of the following is an equation of the line perpendicular to 3x+6y=12 and passing through (4,0)?
Answer Choices:
A) y=-1/2x+2
B) y=1/2x-2
C) y=-2x+8
D) y=2x-8

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

3x+6y=12
6y=-3x+12
x=-3x/6+12/6
x=-x/2+2 this line has a slope of -1/2 therefore a line perpendicular to this line has a slope of it's negative reciprical or 2.
thus the formular for the unknown line is
0=2*4+b
b=-8 or
y=2x-8


Answer by stanbon(75887)

Which of the following is an equation of the line perpendicular to 3x+6y=12 and passing through (4,0)?
----------------------
The given line has a slope of m=-3/6=-1/2
The line perpendicular to the given line must
have a slope of +2/1=2
-------------------
Only y=2x-8 has the correct slope.
-------
Cheers,
Stan H.